# Systematic assessment of various universal machine-learning interatomic potentials

### v1

@@ 2.2 Atomic and lattice relaxations
To test the accuracy of the different uMLIPs further, we prepare a dataset with 19998 materials consisting of unary and binary (with 6903 element combinations) phases in MP. It is worthwhile to highlight that the MP database contains structures relaxed with VASP 44 with the PBE functional. The distribution of the chosen dataset among the chemical elements is reported in Fig. 2. We use the different uMLIPs to perform energy calculations both with no structural optimization (we call these “one-shot” calculations) and with structural optimization. The latter are of two types: (i) only the atomic positions are relaxed (we call these “ion-relax” calculations), and (ii) both atomic positions and cell parameters (i.e., lattice parameters and angles) are relaxed (we call these “cell-relax” calculations). For the structural optimizations, in some cases the calculation stops due to errors while building the graph representation (e.g., given that isolated atoms are found in the structure). Calculations for which this problem appears are tagged as problematic. In another non-negligible number of cases, that we tag as unconverged, the relaxation algorithm is not able to reach the stopping criterion before 150 steps. Finally, we tag as converged the calculations where the stopping criterion is met in less than 150 steps. Additional details on the relaxation algorithm are given in Sec. 4.
@@
In Fig 3, we report the number of occurrences of the problematic, unconverged, and converged tags from the different uMLIPs for the two types of structural optimization. When only the atomic positions are relaxed (Fig 3(a)), the fraction of problematic or unconverged cases is very limited. The best performance is achieved by MatGL (with only 0.2% of unconverged cases), while the worst case is M3GNet (with 2.4% of problematic cases). The fraction of unconverged calculations increases significantly (from 3 to more than 200 times across the uMLIPs) when both atomic positions and cell parameters are optimized (Fig 3(b)). M3GNet and MatGL (both about 60% of unconverged cases) are the most critical cases. The high percentage of unconverged calculations from these two uMLIPs is probably related to the large values of the stress tensors along the structural optimization, which prevent the cell to relax. Full relaxation with CHGNet, MACE, and ALIGNN lead to significantly lower fractions of unconverged results (1.2%, 1.2%, and 4.4%, respectively), with CHGNet and MACE performing very robustly. In the following, we discuss only results from the converged geometry optimizations.
@@
To compare the predictive performance of these uMLIPs regarding energy levels, we also used a benchmark based on the formation energy, defined as
@@
where is the total energy for the phase of interest, and reflect the composition of the compound with their respective fractions and , and and the lowest possible energies of the elemental compounds and , respectively. The ability of the uMLIPs in predicting formation energies is evaluated again as a difference from the MP values, i.e.,
with and being the formation energy from Eq. (3) calculated with MP and the uMLIP model, respectively. M3GNet had to be excluded due to the problematic calculations. Indeed, we could not obtain any energy for a number of elemental compounds (e.g., K, and Rb). Therefore, the number of available formation energies would have been significantly lower for M3GNET than for other uMLIPs. For the same reason (i.e., maximizing the number of formation energies in our dataset), we chose to work with the one-shot results.
@@
In Fig. 4 we report the distribution of (Eq.(4)) for the one-shot calculations using the different uMLIPs. Figure 5 shows the distribution of the absolute mean values of on the chemical space for the different uMLIPs, i.e., the average of the absolute value of over all the chemical systems containing a given element. CHGNet has the best performance for chemical systems containing the transition metals V, Cr, and W, which is likely due to the inclusion of the magnetic moments in this model. MatGL and MACE are less performant for O, V, Fe, and W, where CHGNet and ALIGNN perform very well instead. Besides, MatGL and MACE have slightly poorer performance with respect to CHGNet for chemical compositions including halogens. The performance of the different uMLIPs for (Eq. (4)) for the one-shot calculations is reported in Table 1. CHGNet shows better performance than the other uMLIPs, with the smallest MAE and RMSE and the highest R2, which is 0.081, 0.158 and 0.974, respectively. In turn, CHGNet outperforms the other uMLIPs in the one-shot calculations (Fig. 4 and Table 1), while having almost the same number of converged results (i.e., successful structural optimizations) as MACE, and as double as the number of converged results as MatGL (Fig. 3).
@@
In Fig.6 and Table 2 we summarize the ability of the different uMLIPs (cell-relax calculations) to predict cell parameters, i.e., lattice parameters, angles, and volume. As done for the formation energies (Fig. 4, Eq. (4), and Table 1), we compare the uMLIPs results with the MP values, e.g., for the volume we consider
@@
where and are the cell volume from MP and from the uMLIP model (cell-relax calculations), respectively. In Fig.6 we report the distribution of . CHGNet and ALIGNN outperform the other uMLIPs, with a narrower distribution of . M3GNet is the worst performing, while MatGL and MACE have a somehow intermediate performance. In Table 2, we also report the Mean Absolute Relative Error (MARE) on the predicted lattice parameters (lengths and angles). We observe that CHGNet and MACE show similar performance. In fact, the violin plots are strongly influenced by the presence of outliers. ALIGNN also has a very good performance, while M3GNet and MatGL present significant MARE values (not to mention that they also lead to the highest numbers of unconverged cases).
@@
| CHGNet | 0.081 | 0.158 | 0.974 |
| MatGL | 0.158 | 0.319 | 0.892 |
| MACE | 0.148 | 0.308 | 0.900 |
| ALIGNN | 0.129 | 0.249 | 0.939 |
| uMLIP | a | b | c | |||
| CHGNet | 2.3 | 2.3 | 2.7 | 0.9 | 0.7 | 1.3 |
| M3GNet | 33.4 | 33.3 | 33.5 | 14.3 | 14.2 | 14.1 |
| MatGL | 34.0 | 33.5 | 34.3 | 14.5 | 14.4 | 14.4 |
| MACE | 3.0 | 3.2 | 3.3 | 0.9 | 0.7 | 1.3 |
| ALIGNN | 6.9 | 6.5 | 7.1 | 1.8 | 1.7 | 2.1 |
@@ 2.3 Vibrational properties
In this section, we analyze the capability of uMLIPs to reproduce the vibrational properties of crystalline materials using accurate ab initio results reported in a previous work 39 as a reference. From the structures in Ref. 39 we select those whose energy above hull is zero both in the MP database and from the uMLIP calculations (cell-relax), leading to 101 structures. Phonons are computed using the finite displacement method as implemented in the phonopy package 50, 51. Further details on the protocol employed to compute phonons with uMLIPs are provided in Section 4. By exploiting the finite displacement method to compute phonons, this study indirectly probes the quality of the uMLIPs-calculated forces when atoms are slightly displaced away from the equilibrium positions. From a methodological point of view we note that, already when ab initio engines are used, the accuracy of phonon calculations from the finite displacement method is rather sensitive to the quality of the force calculations in the supercell. Since by construction these forces are less accurate when calculated from uMLIP models than from ab initio methods, it is reasonable to expect some non-negligible discrepancy between uMLIPs and ab initio phonon calculations. Also, as discussed in more detail in Sec. 4, the MLIP architectures at the basis of the universal models considered in this study cannot predict the long-range dipolar contributions to the interatomic force constants of polar materials 52, 53. In order to obtain a reliable Fourier phonon interpolation in the case of polar materials, the uMLIP-calculated forces should therefore be augmented with the ML electronic dielectric tensor and the Born effective charges which are not available at present (see Sec. 4 for further details).
For these reasons, when comparing uMLIPs results with ab initio data, we choose a rather generous metrics, i.e., the MAE between the uMLIPs and the ab initio phonon band structures computed along a high-symmetry -path 54:
In Eq. (6), is the phonon energy in meV, is the branch index, is the number of wavevectors used to sample the -path, and , are the phonon energies computed using the uMLIP model and density-functional perturbation theory (DFPT) with ABINIT 39, respectively. Table 3 reports the minimum, maximum and average values of the MAE in the phonon band structures obtained from the different uMLIPs (Eq. (6), while the MAE distribution in terms of violin plots is given in panel (a) of Fig. 7.
@@
| CHGNet | 0.818 | 16.318 | 7.711 |
| M3GNet | 2.033 | 15.848 | 6.925 |
| MatGL | 2.678 | 26.299 | 7.696 |
| MACE | 0.312 | 5.573 | 3.682 |
| ALIGNN | 11.792 | 49.335 | 30.018 |
However, from panel (a) of Fig. 7 and Table 3, as anticipated above, we observe a non-ideal agreement between uMLIPs-calculated and ab initio-calculated vibrational properties. In addition to the fundamental reasons discussed already, a source of discrepancy could come, in this particular study, from the fact that the PBEsol exchange-correlation functional was used in Ref. 39 while the uMLIPs have been trained using PBE data with the inclusion of U-corrections for certain systems (the only exception being ALIGNN that was trained on optB88vdW data). In panel (b) of Fig. 7 we report the phonon band structure for the system with the lowest MAE (0.312 meV, obtained from MACE, Table 3), which is SrBr (mp-567744). Conversely, in panel (c) of Fig. 7 we report the phonon band structure for the system with the highest MAE (49.335 meV, obtained from ALIGNN, Table 3), which is BeS (mp-422). For the latter, the uMLIP predicts vibrational instabilities that are not observed in the ab initio results.
To summarize, our results indicate that presently available uMLIPs can predict ab initio vibrational properties with a typical error of 3.682 meV in the best case scenario. This value should be considered as a lower bound as the discrepancy is expected to increase significantly if the electronic dielectric tensor and the Born effective charges cannot be predicted with ML techniques.
In the opinion of the authors, the predictive behavior of uMLIPs for vibrational properties can be improved by training new uMLIPs with larger weights for the forces loss function but it is also clear that for accurate ML-based predictions in polar materials, one needs uMLIPs capable of inferring the long-range part of the dynamical matrix. All this being said, we believe that uMLIPs represent an efficient and promising approach to perform an initial screening for vibrational and thermodynamic properties, especially in a high-throughput context in which high accuracy is not necessarily needed.
@@ 3 Conclusions
We present a systematic assessment of various universal machine-learning interatomic potentials (uMLIPs) by investigating their capability to reproduce ab initio results for several important physical properties such as equation of states, relaxed geometries, formation energies and vibrational properties. Among the considered uMLIPs, we find that CHGNet outperforms the others, showing superior precision in predicting the relaxed geometry and formation energy, also considering the number of converged results. CHGNet also has some advantages when considering systems containing elements such as O and some of the transition metal elements. MACE, on the other hand, excels when predicting vibrational properties and, at the same time, is able to infer all the other physical properties with good precision. For what concerns ALIGNN, despite trained with different dataset, it still shows excellent results when predicting lattice parameters and formation energies. M3GNet and MatGL reveal to be the most problematic ones when performing geometry optimization at variable cell. This underscores the need for further optimization and training to fully exploit the capability of ML techniques across a broader range of applications. The choice of a particular uMLIP for specific applications should take into account an appropriate balance between accuracy and computational efficiency. Future work should aim at enhancing the performance of these potentials further, particularly focusing on areas where current uMLIPs exhibit limitations such as more accurate prediction of forces and stresses or the capability of learning Born effective charges and electronic dielectric tensors that are crucial for the vibrational properties of polar materials. Our work will hopefully pave the way towards a more systematic assessment of uMLIPs in different scenarios and the establishment of a standardized benchmark set that can be used to gauge the precision and transferability of uMLIPs.
@@ 4 Methods
The calculations are performed with the AbiPy package 40, 41, more specifically the abiml.py script that provides a unified interface that allows one to perform different types of calculations such as structural relaxations, molecular dynamics or NEB using the algorithms implemented in ASE 55 and different uMLIPs as calculators. The following versions are used to produce the results reported in this work: Python 3.11, Pymatgen 2023.7.17, AbiPy 0.9.6, Abinit 9.8.4, M3GNet:0.2.4, CHGNet:0.3.2, ALIGNN: 2023.10.1, MatGL: 0.9.1, MACE: mace_mp_0 with mace-torch 0.3.4, and ASE: 3.22.1.
Structural relaxations are performed using the ASE optimizer, employing the BFGS algorithm. To ensure good trade-off between accuracy and computational cost, the stopping criterion fmax is set to 0.1 eV/Å. If the stopping criterion is not met at a maximum number of iterations set to 150, the calculation is considered unconverged. This is justified by the fact that all the initial structures in our dataset are taken from the Materials Project 19, where they have been already relaxed with VASP with the PBE functional, and are supposed to be included in the set used to train the uMLIPs (with the exception of ALIGNN).
For the phonon calculations, we first perform a structural relaxation with the uMLIPs from the crystalline structures employed in 39. Only the atomic positions are relaxed at fixed cell in order to avoid spurious effects due to the change of the lattice parameters with respect to the reference DFT results.
The relaxed configuration is then used to compute vibrational properties
using the finite displacement method implemented in phonopy 50, 51
with a displacement of 0.01 Å.
In each calculation, the real-space supercell is matched with the -mesh employed in 39 to compute the dynamical matrix with the DFPT part of Abinit 40, 41.
It is noteworthy that the uMLIPs employed in this work lack the capability to predict the electronic dielectric tensor and the Born effective charge tensor where is the index of the atom in the unit cell.
As discussed in 52, 53, these quantities are needed to
model the long-range dipolar contribution to the interatomic force constants in polar materials.
This term is indeed responsible for the LO-TO splitting as well as for the non-analytical behaviour of the vibrational spectrum near .
Its correct numerical treatment is therefore crucial to obtain
an accurate Fourier interpolation of the dynamical matrix at arbitrary -points.
@@
and obtained with Abinit to model long-range interactions.
The ab initio phonon dispersions is obtained by using the anaddb post-processing tool using the DDB files retrieved from the MP database to compare the different phonon dispersions between uMLIPs and ab initio DFPT results.
Figure 1 and
the heatmaps of the periodic
table in the Supplemental Material
have been produced
using the ab initio results and the open source python scripts available in
the acwf-verification-scripts github repository.

### v3

@@ 2.3 Structural optimization and formation energies
To test the accuracy of the different uMLIPs further, we prepare a dataset with 19998 materials consisting of unary and binary (with 6903 element combinations) phases in MP. It is worthwhile to highlight that the MP database contains structures relaxed with VASP 52 with the PBE functional 25. The distribution of the chosen dataset on the chemical elements is reported in Fig. 2. We use the different uMLIPs to perform energy calculations both with no structural optimization (we call these “one-shot” calculations) and with structural optimization. The latter are of two types: (i) only the atomic positions are relaxed (we call these “ion-relax” calculations), and (ii) both atomic positions and cell parameters (i.e., lattice parameters and angles) are relaxed (we call these “cell-relax” calculations). For the structural optimizations, in some cases the calculation stops due to errors while building the graph representation (e.g., given that isolated atoms are found in the structure). Calculations for which this problem appears are tagged as problematic. In another non-negligible number of cases, that we tag as unconverged, the relaxation algorithm is not able to reach the stopping criterion before 150 steps. Finally, we tag as converged the calculations where the stopping criterion is met in less than 150 steps. Additional details on the relaxation algorithm are given in Sec. 4.
@@
In Fig 3, we report the number of occurrences of the problematic, unconverged, and converged tags from the different uMLIPs for the two types of structural optimization. When only the atomic positions are relaxed (Fig 3(a)), the fraction of unconverged cases is very limited (about 0.30.4% of the total for all the uMLIPs), while only ALIGNN presents some problematic cases worth to be noted, but still very small (0.2%). The fraction of unconverged calculations increases significantly when both atomic positions and cell parameters are optimized (Fig 3(b)), especially for M3GNet (2.5%) and ALIGNN (4.4%). CHGNet and MACE perform more robustly, with lower fractions of unconverged results (both 1.2%). In the following, we discuss only results from the converged geometry optimizations.
@@
To compare the predictive performance of these uMLIPs on the energetics, we use as a benchmark the formation energy, defined as
@@
where is the total energy per atom of the phase of interest with , the constituent elements, and are the fractions of and , respectively, and and are the lowest possible energies per atom of and , respectively. The ability of a uMLIP in predicting formation energies is evaluated as a difference from the MP values, i.e.,
with and being the formation energy (Eq. (3)) from the MP and calculated with the uMLIP model, respectively.
We choose to work with the one-shot energies as this allows to avoid spurious effects due to the different relaxed geometries predicted by the uMLIPs. It is important to note that all our calculations incorporate energy corrections. For CHGNet and M3GNet, these corrections are already encoded in the uMLIPs, as both models were trained using corrected energies 61, 62, 63. Conversely, for MACE, the corrections are applied as a post-processing step (see Sec. 4 and Refs. 81, 82 for details) since this model was trained on uncorrected entries 33. ALIGNN, on the other hand, was trained using the JARVIS-DFT dataset with OptB88vdW functional 32, and does not use any corrected energies, but rather direct DFT outputs. In this case, we compute Eq.(4) with and without energy corrections, and we select the best-agreement set of results.
@@
Figure 4 shows the distribution of the absolute mean values of on the chemical space for the different uMLIPs, i.e., the average of the absolute value of over all the chemical systems containing a given element. CHGNet and MACE have the best performance over all the periodic table. For what concerns the elements in the transition metal series, CHGNet and MACE perform significantly better than M3GNet and ALIGNN, especially for elements like V, Mo, and W. In the case of CHGNet, this might be related to the inclusion of the magnetic moments in the training process. CHGNet and MACE outperform M3GNet and ALIGNN also for what concerns the calchogens (O, S, Se, Te) and some halogens (F, Cl, I). ALIGNN is also less performant for actinide elements Np and Pu compared to other uMLIPs. In Fig. 5, we report the distribution of (Eq.(4)) for the one-shot calculations using the different uMLIPs, while the performance of each model is summarized in Table 2. MACE shows better performance than the other uMLIPs, with the smallest MAE and RMSE and the highest R2 (0.044, 0.101, and 0.989, respectively). In turn, MACE outperforms the other uMLIPs in the one-shot calculations (Fig. 5 and Table 2), while displaying, together with CHGNet, the highest number of successful structural optimizations (or converged results), i.e., 88.8%, while M3GNet and ALIGNN display 87.5% and 85.6%, respectively (Fig. 3).
@@
To examine the ability of the different uMLIPs to predict the cell geometry (i.e., lattice parameters, angles, and volume) in the cell-relax calculations, we compare the uMLIPs results with the MP values, as done for the formation energy (Eq. (4)). E.g., for the volume we consider the relative difference
@@
where and are the cell volume from MP and from the uMLIP model cell-relax calculations, respectively. The distribution of is reported in Fig. 6. CHGNet and M3GNet outperform the other uMLIPs, with a narrower distribution of . ALIGNN performs the worst, while MACE has an intermediate performance. In Table 3, we report the Mean Absolute Relative Error (MARE) on the predicted lattice parameters, angles, and volumes. We observe that MACE shows similar performance with respect to CHGNet and M3GNet for , , and , , , but its performance on the volume is significantly worse than the other two uMLIPs, as also shown in Fig. 6. One possible origin of this behavior might be the compensation of the errors of the cell parameters in the case of M3GNet and CHGNet, which would not occur in MACE. Another reason could be the presence of more outliers (extreme large deviations) in the case of MACE, as documented from the histograms of the relative volume differences and lattice parameters differences provided in the Supplemental Material.
A similar argument can be applied to the slightly higher MARE on the volume for CHGNet as compared to M3GNet (Table 3). ALIGNN shows larger errors compared to the other uMLIPs (Table 3). To further test the ability of uMLIPs to treat multi-element compounds, we perform ‘one-shot’ and ‘cell-relax’ calculations on 100 randomly chosen quinary materials. There are 4 unconverged cases (4%) for CHGNet and M3GNet, while 2 unconverged cases (2%) for MACE and ALIGNN. This seems to indicate that these uMLIPs can still be used for structural relaxations even for quinaries although additional investigations would be needed to assess the transferability in this part of the chemical space. In the Supplemental Material we report the comparison of formation energies and cell geometries of quinaries with the respective MP reference data, as done for the unary and binary compositions (Figs. 5 and 6). This comparison indicates that MACE still gives the most reliable prediction on formation energies, while CHGNet shows the best performance for cell geometry predictions.
@@
| CHGNet | 0.054 | 0.105 | 0.988 |
| M3GNet | 0.172 | 0.316 | 0.896 |
| MACE | 0.044 | 0.101 | 0.989 |
| ALIGNN | 0.137 | 0.2226 | 0.947 |
| uMLIP | volume | a | b | c | |||
| CHGNet | 3.16 | 2.03 | 2.07 | 2.44 | 0.75 | 0.62 | 1.19 |
| M3GNet | 2.97 | 2.04 | 2.09 | 2.46 | 0.89 | 0.73 | 1.24 |
| MACE | 5.22 | 2.01 | 2.11 | 2.58 | 0.73 | 0.59 | 1.13 |
| ALIGNN | 7.85 | 3.42 | 3.42 | 3.61 | 0.94 | 0.86 | 1.32 |
@@ 2.4 Vibrational properties
In this section we analyze the capability of uMLIPs to reproduce the vibrational properties of crystalline materials using accurate ab initio results reported in a previous work 46 as a reference. From the structures in Ref. 46 we select those whose energy above hull is zero both in the MP database and from the uMLIP cell-relax calculations, leading to 101 structures. Phonons are computed using the finite displacement method as implemented in the phonopy package 83, 84. Further details on the protocol employed to compute phonons with uMLIPs are provided in Section 4. By exploiting the finite displacement method to compute phonons, this study indirectly probes the quality of the uMLIPs-calculated forces when atoms are slightly displaced away from the equilibrium positions. From a methodological point of view we note that, already when ab initio engines are used, the accuracy of phonon calculations from the finite displacement method is rather sensitive to the quality of the force calculations in the supercell. Since by construction these forces are less accurate when calculated from uMLIP models than from ab initio methods, it is reasonable to expect some non-negligible discrepancy between uMLIPs and ab initio phonon calculations. Also, as discussed in more detail in Sec. 4, the MLIP architectures at the basis of the universal models considered in this study cannot predict the long-range dipolar contributions to the interatomic force constants, needed to obtain a reliable Fourier phonon interpolation in the case of polar materials 85, 86. The uMLIP-calculated forces should therefore be augmented with the ML electronic dielectric tensor and the Born effective charges, which are not available at present, leading us to augment the uMLIPs forces with ab initio 47, 48 values for these quantities (see Sec. 4 for further details).
For the above reasons, when comparing uMLIPs results with ab initio data, we choose a rather generous metrics, i.e., the MAE between the uMLIPs and the ab initio phonon band structures computed along a high-symmetry -path 87:
In Eq. (6), is the phonon energy in meV, is the branch index, is the number of wavevectors used to sample the -path, and , are the phonon energies computed using the uMLIP model and the density-functional perturbation theory (DFPT) with ABINIT 46, respectively. Table 4 reports the minimum, maximum, and average values of the MAE in the phonon band structures obtained from the different uMLIPs (Eq. (6)), while the MAE distribution is given in panel (a) of Fig. 7.
@@
| CHGNet | 0.82 | 37.34 | 8.12 |
| M3GNet | 0.74 | 40.20 | 10.41 |
| MACE | 0.31 | 17.22 | 3.71 |
| ALIGNN | 5.60 | 75.38 | 29.36 |
From panel (a) of Fig. 7 and Table 4, we observe a non-ideal agreement between uMLIPs-calculated and ab initio-calculated vibrational properties. In addition to the fundamental reasons discussed above, in this particular study a source of discrepancy could come from the fact that Ref. 46 exploits the PBEsol exchange-correlation functional while the uMLIPs have been trained using PBE 25 data with the inclusion of U-corrections 61, 62, 63 for certain systems (in the case of CHGNet and M3GNet) or PBE 25 data without the inclusion of U-correction (in the case of MACE) or optB88vdW 32 data (in the case of ALIGNN). In panel (b) of Fig. 7, we report the phonon band structure for the system with the lowest MAE (0.312 meV, obtained from MACE, Table 4), which is Sr4Br8 (mp-567744). Conversely, in panel (c) of Fig. 7, we report the phonon band structure for the system with the highest MAE (75.375 meV, obtained from ALIGNN, Table 4), which is Be2C (mp-1569). For the latter, the uMLIP predicts vibrational instabilities that are not observed in the ab initio results. However, we also note that even for the lowest-MAE material (panel (b) of Fig. 7), the high-energy region of the phonon band structure is still far from being accurate.
To summarize, our results indicate that presently available uMLIPs can predict ab initio vibrational properties with a typical error of 3.71 meV in the best-case scenario. This value should be considered as a lower bound as the discrepancy is expected to increase if the electronic dielectric tensor and the Born effective charges were predicted with ML techniques. In the opinion of the authors, the predictive behavior of uMLIPs for vibrational properties can be improved by training new uMLIPs with larger weights for the forces loss function but it is also clear that for accurate ML-based predictions in polar materials, one needs uMLIPs capable of inferring the long-range part of the dynamical matrix. All this being said, we believe that uMLIPs represent an efficient and promising approach to perform an initial screening for vibrational and thermodynamic properties, especially in a high-throughput context in which high accuracy is not necessarily needed.
@@ 3 Conclusions
We present a systematic assessment of various universal machine-learning interatomic potentials (uMLIPs) by investigating their capability to reproduce ab initio results for several important physical properties such as equation of state, relaxed geometries, formation energies, and vibrational properties of an extensive set of crystalline materials. Among the considered uMLIPs, we find that MACE shows superior accuracy in predicting formation energies and vibrational properties, and CHGNet and M3GNet are outstanding for relaxed geometry predictions. MACE and CHGNet show superior performance on the formation energy prediction along the periodic table, especially when considering systems containing transition metal elements, chalcogens, and halogens (with Br being accurately reproduced only by CHGNet). M3GNet exhibits relatively high errors, leading to a lower R2 value compared to the other uMLIPs in predicting formation energies. Despite this, its ability to accurately predict volume and lattice parameters remains high, and it has intermediate performance in predicting vibrational properties. For what concerns ALIGNN, despite trained with different dataset, it still shows relatively good results when predicting formation energies. However, it also reveals to be the most problematic model when performing geometry optimization at variable cell and for predicting vibrational properties. These results underscore the need for further optimization and training of the currently available uMLIPs to fully exploit the capability of ML techniques across a broader range of applications. The choice of a particular uMLIP for specific applications should take into account an appropriate balance between accuracy and computational efficiency. Future work should aim at enhancing the performance of these potentials further, particularly focusing on areas where current uMLIPs exhibit limitations such as more accurate prediction of forces and stresses or the capability of learning Born effective charges and electronic dielectric tensors that are crucial for the vibrational properties of polar materials. Our work will hopefully pave the way towards a more systematic assessment of uMLIPs in different scenarios and the establishment of a standardized benchmark set that can be used to gauge the precision and transferability of uMLIPs.
@@ 4 Methods
The calculations are performed with the AbiPy package 47, 48, more specifically the abiml.py script providing a unified interface that allows one to perform different types of calculations such as structural relaxations, molecular dynamics, or NEB using the algorithms implemented in ASE 88 and different uMLIPs as calculators. The following package versions are used to produce the results reported in this work. Python: 3.11, Pymatgen: 2023.7.17, AbiPy: 0.9.6, Abinit: 9.8.4, CHGNet: 0.3.2, ALIGNN-FF: alignn 2023.10.1, MatGL: 1.0.0, MACE-MP0: mace-torch 0.3.4, and ASE: 3.22.1.
Structural relaxations are performed using the ASE optimizer, employing the BFGS algorithm. To ensure good trade-off between accuracy and computational cost, the stopping criterion fmax is set to 0.1 eV/Å. If the stopping criterion is not met at a maximum number of iterations set to 150, the calculation is considered unconverged. This is justified by the fact that all the initial structures in our dataset are taken from the Materials Project 24, where they have been already relaxed with VASP 52 with the PBE functional 25, and the same structures are supposed to be included in the set used to train the uMLIPs (with the exception of ALIGNN). It should be pointed out that for each composition the total energy obtained from the Materials Project 24 is obtained with energy corrections 61, 62, 63. Therefore, we apply the same correction according to the methods in compatibility.py from the Pymatgen package 81, 82 to MACE since it was trained with Materials Project uncorrected energies.
For the phonon calculations we first perform a structural relaxation with the
uMLIPs from the crystalline structures employed in 46.
The atomic positions are relaxed at fixed cell parameters in order to avoid
spurious effects due to the change of the lattice parameters with respect to the
reference DFT results.
The relaxed configuration is then used to compute vibrational properties using
the finite displacement method implemented in
phonopy 83, 84 with a displacement
of 0.01 Å.
In each calculation, the real-space supercell is matched with the -mesh
employed in 46 to compute the dynamical matrix with the DFPT
part of Abinit 47, 48.
It is noteworthy that the uMLIPs employed in this work lack the capability to
predict the electronic dielectric tensor and the Born
effective charge tensor where is the index of the
atom in the unit cell.
As discussed in 85, 86, these quantities are needed to
model the long-range dipolar contribution to the interatomic force constants in
polar materials.
This term is indeed responsible for the LO-TO splitting as well as for the
non-analytical behaviour of the vibrational spectrum near .
Its correct numerical treatment is therefore crucial to obtain an accurate
Fourier interpolation of the dynamical matrix at arbitrary -points.
@@
and obtained with Abinit to model
long-range interactions.
The ab initio phonon dispersions is obtained by using the anaddb post-processing tool using the DDB files retrieved from the MP database to
compare the different phonon dispersions between uMLIPs and ab initio DFPT
results.
Figure 1 and the heatmaps of the periodic table in the
Supplemental Material have been produced using the ab initio results and the
open source python scripts available in
the acwf-verification-scripts
github repository.